OpenStudy (anonymous):

3x-7(x-2)=10 with steps to solving please

OpenStudy (anonymous):

-7(x-2) do this part first

OpenStudy (anonymous):

distribute it?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

3x-7x+14=10? and then just simplify and stuff?

OpenStudy (anonymous):

set the equation to zero so subtract the 10 on the right hand side

OpenStudy (anonymous):

x=1?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

np

OpenStudy (anonymous):

can you answer another one please? ive forgotten all of my algebra over the summer

OpenStudy (anonymous):

3(180-y) = 2(90-y)

OpenStudy (anonymous):

just remember pemdas,